OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 100 y = x + 2 Select one: a. 2 and 10 b. 10 and -2 c. 6 and -8 d. 6 and 8

OpenStudy (*insert name here*):

Do you know how to start?

OpenStudy (anonymous):

no.

OpenStudy (ranga):

y = x + 2 Substitute above for y in x^2 + y^2 = 100. Simplify. You will get a quadratic equation in x. Solve for x.

OpenStudy (anonymous):

but i have to get 2 numbers.

OpenStudy (ranga):

A quadratic equation will yield two solutions.

OpenStudy (anonymous):

i dont know how to do it

OpenStudy (ranga):

x^2 + y^2 = 100 y = x + 2 Take y from the second equation and substitute in the first equation: x^2 + (x+2)^2 = 100 x^2 + x^2 + 4x + 4 = 100 2x^2 + 4x + 4 = 100 2 is a factor in each term. So divide throughout by 2: x^2 + 2x + 2 = 50 x^2 + 2x - 48 = 0 x^2 + 8x - 6x - 48 = 0 x(x+8) - 6(x+8) = 0 (x-6)(x+8) = 0 x = 6 or -8

OpenStudy (anonymous):

Thank you so much, can you help me with a different problem?

OpenStudy (ranga):

I can guide you but I do not wish to do the entire problem myself.

OpenStudy (anonymous):

no, but it is hard for me because i have never done it before and im doing it online and they dont give examples but if you guide me that would be great

OpenStudy (ranga):

alright.

OpenStudy (anonymous):

look, it is a graph

OpenStudy (ranga):

Take the top line first. What is its y-intercept?

OpenStudy (anonymous):

it would be -4,2?

OpenStudy (ranga):

The y-intercept is -4. IDK where you are getting the 2 from. Can you find the slope of the top line? slope = rise/run

OpenStudy (anonymous):

becuase it is passing through the 2 but not exactly am i wrong?

OpenStudy (ranga):

That would be the x-intercept. It is passing between 2 and 3 and so it will be 2.5 or 5/2. But the y-intercept is where the line cuts the y-axis and it is -4. Getting back to the slope. What is the slope of the top line?

OpenStudy (ranga):

There may be an easier way to solve this without even finding the slope. The y-intercept of the top line is -4 What is the y-intercept of the bottom line?

OpenStudy (anonymous):

between -5 &-6 but i say -5?

OpenStudy (ranga):

Yes -5 is correct. Remember, y intercept is where the line cuts the y-axis. That means x = 0. So put x = 0 and solve for y in each of the answer choices and see which one gives the y-intercepts for the two lines to be: -4 and -5 respectively.

OpenStudy (anonymous):

so for example it would be ? an solve a. 3(0) + 2y = 8 6(0) + 4y = 20

OpenStudy (ranga):

Yes. If the first part of the first choice does not give a -4 or a -5 move on to the next choice. You don't have to do the second part of the first choice.

OpenStudy (anonymous):

okei thank you is there any way we can stay in contact?

OpenStudy (ranga):

What was your answer?

OpenStudy (anonymous):

im working on it

OpenStudy (ranga):

I volunteer here when I have some time. No fixed schedule. But there are many individuals who can help. Right now the tagging, messaging, notifications, etc. have not been working properly for the past week or so. You can tag my name when you post a problem and if the notification works and I am logged in I can visit your page.

OpenStudy (anonymous):

i got B

OpenStudy (ranga):

B) has 3x - 2y = 8 and 6x + 4y = 20. Put x = 0 in each and solve for y. 0 - 2y = 8. y = -4 4y = 20, y = 5 Gives -4 and 5 as y-intercepts. But we want -4 and -5. So it is not B.

OpenStudy (anonymous):

D

OpenStudy (ranga):

Yes!

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

have time for another one?

OpenStudy (ranga):

go ahead. one more.

OpenStudy (ranga):

Plot the three points given under the g(x) table. IF any of the points falls on the parabola f(x) then it means g(x) and f(x) will intersect. If no points of g(x) falls on the parabola it means they won't intersect.

OpenStudy (anonymous):

and how do you do that eample';$

OpenStudy (ranga):

You are given 3 points belonging to g(x) in the table. Treat the x value as the x-coordinate and the g(x) value as the y-coordinate. The first point is (-1, 0). Plot that point on the graph and see if it lies on the parabola. Try the other two points also.